OpenStudy (anonymous):

add 16sqrt27 + 2sqrt27

OpenStudy (anonymous):

18sqrt27

jhonyy9 (jhonyy9):

but this is equal again 18*3sqrt3 =54sqrt3

jhonyy9 (jhonyy9):

because sqrt27=sqrt(9*3)=3sqrt3

OpenStudy (anonymous):

$$ 54 \sqrt{3}$$